A cabin cruiser travels 20 miles in the same time that a power boat travels 40 miles. The cruiser travels 5 mph slower than the power boat. Find the speed of each boat.

Answer by checkley71(8403)

SEEING AS THE TIMES ARE EQUAL THEN D/(R-5) FOR THE CRUISE=D/R FOR THE POWER BOAT:
20/(R-5)=40/R CROSS MULTPLY
40(R-5)=20R
40R-200=20R
40R-20R=200
20R=200
R=200/20
R=10 MPH FOR THE POWER BOAT.
10-5=5 MPH FOR THE CRUISER.
PROOF
20/5=40/10
4=4